A box contains 6 black and 5 red balls. Two balls are drawn one after another from the box without replacement. The probability for both balls to be red is
- A. 3/11
- B. 2/11
- C. 5/11
- D. 25/121
Answer: B. 2/11
Explanation: P(R1) = 5/11. P(R2|R1) = 4/10. P(Both) = 5/11 * 4/10 = 20/110 = 2/11.